OpenStudy (anonymous):

What are the zeros of the polynomial function: f(x) = x3 + x2 – 6x ?

OpenStudy (anonymous):

0,2.-3 my answer

OpenStudy (anonymous):

\[x^3 + x^2 – 6x =0\] \[x(x^2+x-6)=0\] \[x(x+3)(x-2)=0\] looks like you are right!